# Enhancing Efficiency in Multidevice Federated Learning through Data Selection

### v4

@@ 3.2. Model Initialization
To partition a DNN into an encoder $E$ and a classifier $C$, one option is to use pre-trained benchmark DNNs; considering numerous well-trained models, e.g., in Model Zoo (Koh, 2019). Usually, DNN architectures from a candidate space of pre-trained models have already been well developed, with efficient quantization and compression capabilities, to be deployed on UCDs for inference (Fedorov et al., 2019; Banbury et al., 2021; Xu et al., 2022; Liberis et al., 2021; Malekzadeh et al., 2021; Li et al., 2021a). We aim to utilize the full capability of UCDs for on-device training of the classifier $C$. Thus, we need to design the classifier such that it can run on the limited memory and computation power available on these devices.
To achieve these, we can iteratively look into candidate architectures that fit in the limited memory available on these UCDs. A robust mechanism in this direction is to obtain a history of memory usage of a particular device which reflects the typical memory availability on the UCD. Formally speaking, the iterative process can be viewed as an optimization problem described as
(1) $C=\min\{m_{1},m_{2},\ldots,m_{k}|\mathbb{P}(m_{k})\times\mathbb{B}<\mathbb{M}\},$
where $\mathbb{P}(m_{k})$ provides the number of parameters of the candidate architecture $m_{k}$, $\mathbb{B}$ is the number of bytes used to store the parameters and intermediate results in computations, and $\mathbb{M}$ is the memory available in the device for FL participation. The parameter $\mathbb{M}$ can be measured and logged as empirical observations of the UCD memory usage, to be sent to the server as we discussed in Step \⃝raisebox{-0.9pt}{1} in Figure 4.
In this paper, we limit our classifier architectures with fully connected layers; thus, we find a solution by either reducing the number of neurons or the layers to find a potential architecture for $C$ that can fit into the requirements for $\mathbb{M}$ in Equation (1). In the same vein, we tackle the problem of low computing resources by choosing the smallest architecture for $C$ and then deploying it on the UCDs.
@@ 3.3. Data Selection
All available data $D$ at round $r$ are fed into the model $M$ to compute their loss values $\ell_{r}(f_{M}(\mathbf{X}),\mathbf{y})$. We consider lower loss values as an indication of being less important, and higher loss values show the importance of the data (Jiang et al., 2019).
@@
(2) $\begin{cases}\mathcal{P}_{N}(\ell_{r+1})=1-[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\alpha}\\ \mathcal{P}_{M}(\ell_{r+1})=[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\beta}\end{cases}$
where $\alpha$ and $\beta$ parameters determine the level of selectivity. At round $r+1$, for a sample with loss value $\ell_{r+1}$, the sample is selected into $D_{N}$ with probability of $\mathcal{P}_{N}(\ell_{r+1})$. This implies that the lower the loss value for a sample, the greater the probability it will be discarded. Similarly, $\mathcal{P}_{M}$ is used to select samples with the highest losses as $D_{M}$; which means that the higher the loss value for a sample, the greater the probability it will be chosen for training the entire model.
@@
(3) $\mathcal{P}_{M}^{+}(\|g_{r+1}\|)=[\text{CDF}^{g}_{r}(\|g_{r+1}\|)]^{\gamma}$
321 Inputs: (1) $\mathcal{A}$: the set of $A$ clients each having two devices, one UCD and one AP, (2) $D^{k}$: local dataset for each client $a\in\{1,\dots,A\}$, (2) $f_{E}$: the encoder part with parameters $W_{E}$, (3) $f_{C}$: the classifier part with parameters $W_{C}$, (4) $R$: the total number of FL rounds. for
$r\in\{1,\dots,\nicefrac{{R}}{{2}}\}$ rounds
 do
6      54 $K$ clients $\Leftarrow$ randomly select $K$ clients from a total of $A$, using uniform sampling — Training on UCDs — for $k\in\{1,...,K\}$ do
11            10987 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D^{k}$ (all the samples in the local dataset of client $k$ UCD) $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values) $D_{N}^{k},D_{C}^{k},D_{M}^{k}\Leftarrow$ based on $\ell^{k}$, perform data selection as detailed in Section 3.3 for all $(\mathbf{X}^{k}_{j},\mathbf{y}^{k}_{j})\in D_{C}^{k}$ do
16                  15141312 $g^{k}_{C}={\partial\ell^{k}}/{\partial W_{C}}$ (compute $f_{C}$’s gradients on $D_{C}^{k}$ to update $f_{C}$)$W_{C}^{k}=Optimizer(W_{C}^{k},g^{k}_{C})$ (update $f_{C}$’s parameters using gradients) $D^{+}\Leftarrow$ based on $g^{k}_{C}$, perform data selection as detailed in Section 3.3 $D_{M}^{k}=D_{M}^{k}\cup D^{+}$ (enhance $D_{M}^{k}$ by high-value gradients data points)
20      191817 — Server Side Aggregation on the Classifier Part — $W_{C}\leftarrow\frac{1}{K}\sum_{k\in K}W_{C}^{k}$ $W_{C}^{k}\leftarrow W_{C}$ (update all clients with new $W_{C}$) — Training on APs — for $k\in\{1,\dots,K\}$ do
25            24232221 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D_{M}^{k}$ (all selected samples for training $f_{E}$ on client $k$ AP) $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values) $g^{k}={\partial\ell^{k}}/{\partial(W_{E},W_{C})}$ (compute gradients on $D_{M}^{k}$ to update $f_{E}$ and $f_{C}$) $W_{E}^{k},W_{C}^{k}=Optimizer(W_{E}^{k},W_{C}^{k},g^{k})$ (update all parameters using gradients)
28      2726 — Server Side Aggregation on both Encoder and Classifier Parts — $\{W_{E},W_{C}\}\leftarrow\frac{1}{K}\sum_{k\in K}\{W_{E}^{k},W_{C}^{k}\}$$\{W_{E}^{k},W_{C}^{k}\}\leftarrow\{W_{E},W_{C}\}$ (update all clients with the new aggregated model)
29 return $\{W_{E},W_{C}\}$
Algorithm 1 Multidevice FL with Data Selection and Partition-based Training
3.4. Partition-based Training and Aggregation
The training process is conducted on both UCDs and APs. While UCDs only train the classifier $C$ on data $D_{C}$, APs train the complete model $M$ on data $D_{M}$. Such procedure allows the model to be (partially) updated when UCDs are offline, in addition to the full updates when UCDs have a connection to the internet. This better utilizes UCDs’ spatiotemporal richness and consequently improves the training efficiency in our design requirements.
Algorithm 1 shows the process of partition-based training and aggregation in Centaur. We set the algorithm to run $\nicefrac{{R}}{{2}}$ rounds, considering that UCD training and AP training proceed iteratively in one round. All training is conducted on the client’s devices either on UCDs or APs. Then with all updated weights, the server performs Federated Averaging (FedAvg) (McMahan et al., 2017b) to obtain one aggregated global model. The new global model needs to be distributed to all clients UCDs and APs such that (i) the training on AP has an updated classifier, and (ii) the training on UCD in the next round has both an updated encoder and classifier.
4. Metrics
@@
To evaluate the performance of our FL framework, compared to other FL alternatives, we use the following metrics:
@@
2) Multiply–Accumulate (MAC). This operation is a common step that computes the product of two numbers and adds that product to an accumulator ($a\leftarrow a+(b\times c)$); a fundamental operation for both any DNN layers during training and inference. We use fvcore (FAIR, 2019) library to compute the number of MAC operations. Since fvcore only supports counting MACs in a forward pass, and to count MACs in a backward pass, we use the heuristic that FLOPs (i.e., double MACs) ratio of the backward-forward pass is typically between 1$\times$ and 3$\times$ and most often is 2$\times$ based on models’ specific layer types, according to previous observations (Marius Hobbhahn, 2021).
3) Bandwidth. As the model size (i.e., the encoder or the classifier) can be different in each FL round, we use fvcore to count the number of parameters that are communicated in each round. Based on the model size and number of communications among UCDs, APs, and the server, we compute the amount of bandwidth that is consumed. When it is a case, we also count the number of sample points that are uploaded from UCDs to APs.
4) Latency. The latency usually has a linear relation with MAC operations; due to the lack of specialized accelerators (Liberis et al., 2021; Lin et al., 2022). Thus, we estimate the latency of model training based on the processor’s frequency:
(4) $\text{Latency}=c*\frac{\text{Total MACs}}{\text{Processor Frequency}}.$
The ratio between MAC operations and the processor’s instructions, $c$, is typically between $1$ and $2$ based on specific instruction sets/compilers. For simplicity, we assume that each MAC operation translates to two instructions in an MCU (i.e., $c=2$). Notice that existing processors (e.g., Intel’s Load Effective Address) completes one MAC in one instruction (WiKiBooks, 2021)), and this only scales the experimental results and does not change the final conclusions. Finally, the latency of communication is estimated based on the total amount of data needed to be transmitted, divided by the up-link speed or the down-link speed of devices.
5) Energy. This is the total execution time multiplied by the processor’s consumed power per unit of time. The energy consumption of communication can also be calculated based on the total time of transmitting data multiplied by the transmitters’ power per unit of time. We set the values for UCDs and APs in our simulations as follows:
(A) As computation resources, we set up:
Device
CPU Frequency (MHz)
Storage Capacity (MB)
Power Capacity (mW/MHz)
Timeout (s)
UCD 100 5 0.05 Inf
AP 2000 4096 1.5 Inf
(B) For communication resources, we set up:
Device
Uplink(Mbit/s)
Downlink (Mbit/s)
Communication Energy (W)
Disconnect Probability
UCD 2 2 0.0001 0.5
AP 10 100 10 0
5. Experimental Setup
Datasets. We use six commonly used datasets. CIFAR10 and CIFAR100 (Krizhevsky et al., 2009) that both include 50K training samples and 10K test samples, with 10 and 100 classes respectively. EMNIST (Cohen et al., 2017) that includes 112K training samples with 47 classes. UCIHAR (Anguita et al., 2013) is a widely used dataset of 30 users performing 6 daily activities; data from the accelerometer and gyroscope sensors were collected by a smartphone worn on the waist. Data from 21 users are used for training and that of the other 9 users for testing purposes. MotionSense (Malekzadeh et al., 2019) also includes accelerometer and gyroscope data from 24 users with a smartphone in the pocket of the trousers who performed 6 activities in 15 trials. We use as test data one trial session for each user and as training data the remaining trial sessions (e.g., one trial of “walking” of each user is used as test data and the other two trials are used as training). PAMAP2 (Reiss and Stricker, 2012) dataset contains data of 13 different physical activities, performed by 9 subjects wearing 3 devices and a heart rate monitor. We use the training-test split provided by PersonalizedFL library (Lu and Wang, 2023).
Models. For image classification datasets, we select four DNNs commonly used in mobile/edge-oriented literature: (1) EfficientNet-v2 (Tan and Le, 2021), (2) MobileNet-v3 (Howard et al., 2019), (3) MNASNet (Tan et al., 2019), and (4) ShuffleNet (Ma et al., 2018), with 5.3M, 2.2M, 2.5M, and 1.4M number of parameters respectively. For human-activity recognition datasets, we borrow the ConvNet architecture proposed in (Chang et al., 2020). All these models consist of an encoder made of multi-layer CNNs followed by one classifier.
FL Settings. We use Flower (Beutel et al., 2020), a customizable open-source FL framework (Python v3.7, Ray v1.11, and Torch v1.12). We run all simulations on a server with 80 Intel Xeon(R) E5-2698 CPUs, 8 Tesla V100 GPUs (16GB), and 504GB system RAM. For all simulations, we use 100 communication rounds, where at each round, UCDs and APs successively train the model. Unless specified otherwise, we consider 100 clients (each having one AP and one UCD). We randomly sample 10% of the clients for training in each round in our initial experiments and later show that our results hold when the client participation ratio is gradually increased to 100%. Data ($D$) are partitioned using Latent Dirichlet Allocation (LDA) (Blei et al., 2003; Beutel et al., 2022) without resampling (LDA-Alpha=1000). We refer to prior research (Blei et al., 2003) about details of LDA’s generative process. After such partition, each client owns $\nicefrac{{|D|}}{{100}}$ local samples.
Figure 5. Test accuracy when using different values for parameters $\alpha$, $\beta$, $\gamma$ in data selection.
@@ 5. Experimental Setup
Baselines. (1) AP Training. In this baseline approach we consider UCDs as data collection apparatuses only and therefore, UCDs upload collected and stored data samples to their connected APs, so that the training only happens on APs. In this case, the complete model (both the encoder and classifier) is updated because APs in general are considered to have sufficient resources. (2) UCD Training. UCDs do not upload the collected data samples. Instead, they conduct training on the data locally. However, as UCDs are typically resource-constrained, only the classifier is considered to be trained on-device. This means that the encoder part is considered to be frozen in this setting.
Mobility Model. In our case, the idea of the mobility model signifies how the mobility patterns of the UCDs (or users) impact the connectivity of the UCD to its AP and eventually to the internet. We define an exclusive-Online Association Matrix $\Omega$, which is a binary matrix representing the user’s exclusive location at a given time instance. Mathematically, $\Omega$ is a binary matrix $\mathcal{T}\times\mathcal{S}$ where $\mathcal{T}$ and $\mathcal{S}$ represent the temporal and spatial granularity of $\Omega$ marix, respectively. For example, the rows may represent the time zones of the day (early morning, morning, afternoon, evening, and late night), during which the user moves between three locations like home, office, and a public park, represented by the columns of $\Omega$. Furthermore, we ensure that each row’s sum equals unity, which means the user is present exclusively at a unique location in a given temporal instance. Finally, to simulate the connectivity patterns, we define a connectivity matrix $\lambda$, a single row-vector of dimension $\mathcal{S}$, denoting the connectivity probability across different locations. More specifically, for evaluating Centaur during mobility, we generate the global connectivity matrix $\lambda$ and the generate matrix $\Omega$, both chosen uniformly at random, for each user to simulate the mobility scenario.
6. Evaluation Results
We present the results when running Centaur compared to conventional FL training methods. We also analyze the impact of data/participation heterogeneity and spatiotemporal coverage on training methods.

### v6

@@ 4.1. Model Initialization
To partition a DNN into an encoder $E$ and a classifier $C$, one option is to use pre-trained benchmark DNNs; considering numerous well-trained models. Usually, DNN architectures from a candidate space of pre-trained models have already been well developed, with efficient quantization and compression capabilities, to be deployed on UCDs for inference fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas; li2021hermes. We aim to utilize the full capability of UCDs for on-device training of the classifier $C$. Thus, we need to design the classifier such that it can run on the limited memory and computation power available on these devices. To achieve these, we can iteratively look into candidate architectures that fit in the limited memory available on these UCDs. A robust mechanism in this direction is to obtain a history of memory usage of a particular device, which reflects the typical memory availability on the UCD.
To make the selection process explicit, we consider a set of lightweight architectures $\{m_{k}\}$: a small set of widely used mobile convolutional models (e.g., SqueezeNet and MobileNet-style variants), alongside fully connected networks with varying depth (1–3 layers) and width (16–128 neurons). These architectures are chosen because they span the typical memory–compute trade-off space encountered in edge devices while still supporting on-device gradient updates. Among the feasible candidates that satisfy the device’s resource constraints, we select the smallest model that achieves an accuracy within a small tolerance of the best-performing candidate. This strategy effectively balances efficiency and performance, ensuring that the chosen classifier maximizes on-device feasibility without incurring a significant loss in learning effectiveness.
Formally speaking, the iterative process can be viewed as an optimization problem described as
$C=\operatorname*{argmin}_{\{m_{1},\dots,m_{K}\}}\;\mathbb{P}(m_{k})\,\mathbb{B}\penalty 10000\ \penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ \mathbb{P}(m_{k})\,\mathbb{B}\leq\mathbb{M}\penalty 10000\ \text{and}\penalty 10000\ \text{acc}(m_{k})\geq\tau^{*},$ (1)
[…]
21Inputs: (1) $\mathcal{A}$: the set of AA clients each having two devices, one UCD and one AP, (2) $D^{k}$: local dataset for each client $a\in\{1,\dots,A\}$, (2) $f_{E}$: the encoder part with parameters $W_{E}$, (3) $f_{C}$: the classifier part with parameters $W_{C}$, (4) $R$: the total number of FL rounds.
30.1cm
54 for $r\in\{1,\dots,\nicefrac{{R}}{{2}}\}$ rounds do
7 6 $K$ clients $\Leftarrow$ randomly select $K$ clients from a total of AA, using uniform sampling
8 0.1cm
9  — Training on UCDs —
11 10 for $k\in\{1,...,K\}$ do
13    12 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D^{k}$ (all the samples in the local dataset of client $k$ UCD)
15    14 $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values)
17    16 $D_{N}^{k},D_{C}^{k},D_{M}^{k}\Leftarrow$ based on $\ell^{k}$, perform data selection as detailed in Section 4.2
19    18 for all $(\mathbf{X}^{k}_{j},\mathbf{y}^{k}_{j})\in D_{C}^{k}$ do
21       20 $g^{k}_{C}={\partial\ell^{k}}/{\partial W_{C}}$ (compute $f_{C}$’s gradients on $D_{C}^{k}$ to update $f_{C}$)
23       22 $W_{C}^{k}=Optimizer(W_{C}^{k},g^{k}_{C})$ (update $f_{C}$’s parameters using gradients)
25       24 $D^{+}\Leftarrow$ based on $g^{k}_{C}$, perform data selection as detailed in Section 4.2
27       26 $D_{M}^{k}=D_{M}^{k}\cup D^{+}$ (enhance $D_{M}^{k}$ by high-value gradients data points)
28  — Server Side Aggregation on the Classifier Part —
30 29 $W_{C}\leftarrow\frac{1}{K}\sum_{k\in K}W_{C}^{k}$
32 31 $W_{C}^{k}\leftarrow W_{C}$ (update all clients with new $W_{C}$)
33 0.1cm
34  — Training on APs —
36 35 for $k\in\{1,\dots,K\}$ do
37
39    383pt $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D_{M}^{k}$ (all selected samples for training $f_{E}$ on client $k$ AP)
41    40 $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values)
43    42 $g^{k}={\partial\ell^{k}}/{\partial(W_{E},W_{C})}$ (compute gradients on $D_{M}^{k}$ to update $f_{E}$ and $f_{C}$)
45    44 $W_{E}^{k},W_{C}^{k}=Optimizer(W_{E}^{k},W_{C}^{k},g^{k})$ (update all parameters using gradients)
46  — Server Side Aggregation on both Encoder and Classifier Parts —
48 47 $\{W_{E},W_{C}\}\leftarrow\frac{1}{K}\sum_{k\in K}\{W_{E}^{k},W_{C}^{k}\}$
50 49 $\{W_{E}^{k},W_{C}^{k}\}\leftarrow\{W_{E},W_{C}\}$ (update all clients with the new aggregated model)
51 return $\{W_{E},W_{C}\}$
Algorithm 1 Centaur: Multidevice Federated Learning via Partition-based Training and Data Selection
@@ 4.2. Data Selection
$\begin{cases}\mathcal{P}_{N}(\ell_{r+1})=1-[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\alpha}\\ \mathcal{P}_{M}(\ell_{r+1})=[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\beta}\end{cases}$ (2)
@@
For samples that are not selected through $\mathcal{P}_{N}$ and $\mathcal{P}_{M}$, and to avoid consuming extra resources of UCDs, we can further compute the gradients of the classifier’s last layer. The norm of these gradients gives us a useful hint about the sample’s importance while requiring much less computation than computing all layers’ gradients katharopoulos2018not. The last layer’s norm at round $r$ is produced during training of the classifier by $g_{r}=\frac{\partial\ell_{r}}{\partial W_{r}}$, where $W_{r}$ is the weights of the classifier $C$’s last layer at round $r$. Following the same idea, we derive a CDFg for gradient norm values and build a queue to keep track of the computed norm of gradient values. Samples with larger gradients have larger impacts on the model’s weights; thus, we also keep these samples for training the entire model $M$ as $E$ might learn new “features” from them. Thus, the probability of adding a sample to $D_{M}$ at round $r+1$ is defined as:
$\mathcal{P}_{M}^{+}(\|g_{r+1}\|)=[\text{CDF}^{g}_{r}(\|g_{r+1}\|)]^{\gamma}$ (3)
While our proposed data selection strategy is primarily empirical, it is motivated by the theoretical principles of importance sampling in deep learning. Prior work has shown that sampling data points in proportion to their loss values or gradient magnitudes can approximate the optimal sampling distribution that minimizes the variance of stochastic gradient estimates and thus accelerates convergence katharopoulos2018not; jiang2019accelerating. In particular, high-loss samples tend to carry larger gradient norms and contribute more substantially to the update direction, while samples with small losses or small gradient norms contribute less. Our combination of loss value and last-layer gradient norm therefore aligns with these theoretical insights: it adaptively emphasizes samples expected to produce high-informative gradients, balancing computational cost with learning efficacy. Nonetheless, we acknowledge that our current formulation does not provide a formal convergence guarantee. Developing theoretical guarantees through variance-reduction or convergence-rate analysis is a promising direction for future work.
4.3. Partition-based Training and Aggregation
The results of §3 highlight why we freeze the encoder and only train the classifier on UCDs. Training just the classifier consumed far less memory and CPU, comparable to running an inference pass, and ran much faster than full-model training. Since most UCDs have far less memory than even a RaspberryPi (e.g. a few MB of storage or only a few hundred KB of SRAM), full-model training on UCDs is generally infeasible. By freezing the encoder on UCDs, Centaur aims to reduce the on-device computation and memory requirements, allowing each UCD to train only the compact classifier locally. The more powerful AP (smartphone or router) can then periodically update the encoder on the offloaded samples, so that the overall model can still learn richer features even though each UCD only fine-tunes the classifier part.
Therefore, our training is conducted on both UCDs and APs. While UCDs only train the classifier $C$ on data $D_{C}$, APs train the complete model $M$ on data $D_{M}$. Such a procedure allows the model to be (partially) updated when UCDs are offline, in addition to the full updates when UCDs have a connection to the internet. This better utilizes UCDs’ spatiotemporal richness and consequently improves the training efficiency in our design requirements.
Algorithm 1 shows partition-based training and aggregation in Centaur, running $\nicefrac{{R}}{{2}}$ rounds, considering that UCD training and AP training proceed iteratively in one round. All training is conducted on the client’s devices either on UCDs or APs. Then with all updated weights, the server performs Federated Averaging (FedAvg) floriginal to obtain one aggregated global model. The new global model needs to be distributed to all clients UCDs and APs such that (i) the training on AP has an updated classifier, and (ii) the training on UCD in the next round has both an updated encoder and classifier.
Centaur is the first FL framework to simultaneously combine model partitioning, data selection, and on-device training. Our approach of partitioning the DNN into encoder and classifier, selecting a subset of data on-device, and performing lightweight on-device updates, has not been addressed jointly in earlier work. In contrast, existing hierarchical FL schemes either train full models on each device or only aggregate updates at the AP without splitting the model.
5. Experimental Setup
Datasets. We use six commonly used datasets. CIFAR10 and CIFAR100 krizhevsky2009learning that both include 50K training samples and 10K test samples, with 10 and 100 classes, respectively. EMNIST cohen2017emnist that includes 112K training samples with 47 classes. UCIHAR Anguita2013APD is a widely used dataset of 30 users performing 6 daily activities; data from the accelerometer and gyroscope sensors were collected by a smartphone worn on the waist. Data from 21 users is used for training, and that of the other 9 users for testing purposes. MotionSense malekzadeh2019mobile also includes accelerometer and gyroscope data from 24 users with a smartphone in the pocket of the trousers who performed 6 activities in 15 trials. We use as test data one trial session for each user and as training data the remaining trial sessions (e.g., one trial of “walking” of each user is used as test data and the other two trials are used as training). PAMAP2 reiss2012introducing dataset contains data of 13 different physical activities, performed by 9 subjects wearing 3 devices and a heart rate monitor. We use the training-test split provided by the PersonalizedFL library PersonalizedFL.
Models. For image classification datasets, we select four DNNs commonly used in mobile/edge-oriented literature: (1) EfficientNet-v2 tan2021efficientnetv2, (2) MobileNet-v3 howard2019searching, (3) MNASNet tan2019mnasnet, and (4) ShuffleNet ma2018shufflenet, with 5.3M, 2.2M, 2.5M, and 1.4M number of parameters respectively. For human-activity recognition datasets, we borrow the ConvNet architecture proposed in chang2020systematic. All these models consist of an encoder made of multi-layer CNNs followed by a classifier.
FL Settings. We use Flower beutel2020flower, a customizable open-source FL framework (Python v3.7, Ray v1.11, and Torch v1.12). We run all simulations on a server with 80 Intel Xeon(R) E5-2698 CPUs, 8 Tesla V100 GPUs (16GB), and 504GB system RAM. For all simulations, we use 100 communication rounds, where at each round, UCDs and APs successively train the model. Unless specified otherwise, we consider 100 clients (each having one AP and one UCD). We randomly sample 10% of the clients for training in each round in our initial experiments and later show that our results hold when the client participation ratio is gradually increased to 100%. Data ($D$) are partitioned using Latent Dirichlet Allocation (LDA) blei2003latent; beutel2022flower without resampling (LDA-Alpha=1000). We refer to prior research blei2003latent about details of LDA’s generative process. After such partition, each client owns $\nicefrac{{|D|}}{{100}}$ local samples.
For each client, we randomly select half of the samples and use them for local training of the model. For both APs and UCDs, we consider by default the number of epochs to be 3 with a batch size of 64. The other half of the samples are available with more offline time to simulate the further collection of extra data. APs and UCDs also train more epochs on the extra data while offline. In this setup, we define every unit offline time contributing to one more epoch of training on $20\%$ of these extra local samples.
To set the hyperparameters for Eqs (2)-(3), we run an experiment on CIFAR-10 with MobileNet-V3. We evaluate the trade-off between test accuracy and the workload imposed by data selection on both UCDs and APs. This experiment can be executed by the server as a preparatory step before deploying Centaur, using small representative datasets that approximate the clients’ data distribution. The results in Figure 4 show that performing data selection reduces latency and energy consumption on UCDs by over 80%. Although this approach increases the workload on APs, this is acceptable given their higher computational capacity, since we observe a substantial improvement in model accuracy. In our experiments, we set hyperparameter $\alpha,\beta,\gamma$ as $5,3,0$, respectively, as this configuration represents the fair balance between accuracy and resource efficiency.
@@ 5. Experimental Setup
Baselines. (1) AP Training. We consider UCDs as data collection apparatuses only and therefore, UCDs upload collected and stored data samples to their connected APs, so that the training only happens on APs. In this case, the complete model (both the encoder and classifier) is updated because APs in general are considered to have sufficient resources. (2) UCD Training. UCDs do not upload the collected data samples. Instead, they conduct training on the data locally. However, as UCDs are typically resource-constrained, only the classifier is considered to be trained on-device. This means that the encoder part is considered to be frozen in this setting.
Mobility Model. In our case, the idea of the mobility model signifies how the mobility patterns of the UCDs (or users) impact the connectivity of the UCD to its AP and eventually to the internet. We define an exclusive Online Association Matrix $\Omega$, which is a binary matrix representing the user’s exclusive location at a given time instance. Mathematically, $\Omega$ is a binary matrix $\mathcal{T}\times\mathcal{S}$ where $\mathcal{T}$ and $\mathcal{S}$ represent the temporal and spatial granularity of $\Omega$ marix, respectively. For example, the rows may represent the time zones of the day (early morning, morning, afternoon, evening, and late night), during which the user moves between three locations like home, office, and a public park, represented by the columns of $\Omega$. Furthermore, we ensure that each row’s sum equals unity, which means the user is present exclusively at a unique location in a given temporal instance. Finally, to simulate the connectivity patterns, we define a connectivity matrix $\lambda$, a single row-vector of dimension $\mathcal{S}$, denoting the connectivity probability across different locations. More specifically, for evaluating Centaur during mobility, we generate the global connectivity matrix $\lambda$ and the generate matrix $\Omega$, both chosen uniformly at random, for each user to simulate the mobility scenario.
6. Evaluation Results
We present the results when running Centaur compared to conventional FL training methods and compare the approaches using different metrics. We also analyze the impact of data/participation heterogeneity and spatiotemporal coverage on training methods.
6.1. Metrics
@@
In addition to the accuracy of the trained global model on a held-out test set we evaluate the performance of our FL framework, compared to other FL alternatives, using the following metrics.
@@
3) Bandwidth. As the model size (i.e., the encoder or the classifier) can be different in each FL round, we use fvcore to count the number of parameters that are communicated in each round. Based on the model size and number of communications among UCDs, APs, and the server, we compute the amount of bandwidth that is consumed. Also, we count the number of sample points that are uploaded from UCDs to APs.
4) Latency. The latency usually has a linear relation with MAC operations, due to the lack of specialized accelerators liberis2021munas; lin2022device. Thus, we estimate the latency of model training based on the processor’s frequency as - $\text{Latency}=c*\frac{\text{Total MACs}}{\text{Processor Frequency}}$. The ratio between MAC operations and the processor’s instructions, $c$, is typically between $1$ and $2$ based on specific instruction sets/compilers. For simplicity, we assume that each MAC operation translates to two instructions in an MCU (i.e., $c=2$). Notice that existing processors (e.g., Intel’s Load Effective Address) complete one MAC in one instruction x86assembly), and this only scales the experimental results and does not change the conclusions. Finally, communication latency is estimated based on the total amount of data needed to be transmitted, divided by the up-link speed or the down-link speed of devices.
5) Energy. This is the total execution time multiplied by the processor’s consumed power per unit time. The energy consumption of communication can also be calculated based on the total time of transmitting data multiplied by the transmitters’ power per unit time. We set the values for UCDs and APs in our simulations as follows. We assume a typical AP to have a CPU frequency of 2GHz, storage capacity of 4GB, power capacity of 1.5mW/MHz with an uplink speed of 10Mbit/s, downlink speed 100Mbit/s, and communication energy of 10W. For a typical UCDs, we assume a CPU frequency of 100MHz, storage capacity of 5MB, power capacity of 0.05mW/MHz with an uplink speed of 2Mbit/s, downlink speed of 2Mbit/s, and communication energy of 0.0001W.